let X1=1 and Xn+1=sqrt(3xn) for all n€N show that (Xn) is bounded monotone and converges to 3

1
Expert's answer
2022-02-28T11:43:11-0500

ANSWER.

Using the method of mathematical induction , we will show that "x_{n}<3" for every "n" .

Observe, that "x_{1}=1<3." From the inductive assumption "x_{k}<3" it follows that "x_{k+1}=\sqrt{3x_{k}}<\sqrt{3\cdot3 }=3" . So "x_{n}<3" for every "n" by the mathematical induction.

Since "x_{n}>0" and "x_{n}-3<0" ,then "x_{n}\cdot (x_{n}-3)<0" or "3x_{n}>x_{n}^{2}" . Hence,

"x_{n+1}=\sqrt{3x_{n}}>x_{n}" .So the sequence is increasing. Since "\left \{ x_{n} \right \}" is increasing and bounded above "\lim _{n\rightarrow\infty}x_{n}=a" exists. The function "f(x)=\sqrt{3x}" is continuous, therefore

"a=\lim_{n\rightarrow\infty}x_{n+1}=\lim_{n\rightarrow\infty}\sqrt{3x_{n }}=\sqrt{3\cdot \lim_{n\rightarrow\infty}x_{n }}=\sqrt{3a}" . Thus "a^{2}=3a" .

Since "x_{n}>x_{1}=1" for every "n" we must have "a\ge1" . Therefore "a=3" and "\lim_{n\rightarrow\infty}x_{n}=3" .